OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (5, 5) m = -3 i got y=20?

OpenStudy (anonymous):

let me guess. b = 20?

OpenStudy (anonymous):

y-5=-3(x-5) y=-3x+20